OpenStudy (anonymous):

Please help! Find the exact value for cos 165° using the half-angle identity. I have gotten to cos82.5°=+-sqrt(1-.96)/2 The answer I got was cos82.5°=.366 Can someone double check my work?

OpenStudy (solomonzelman):

cos 165 = -cos 15 = -cos (30/2) = -sqrt [(1+cos30)/2] = -sqrt [(1+sqrt(3)/2)/2] = -sqrt [2+sqrt(3)]/2 Do you see how I did it?

OpenStudy (solomonzelman):

\[-\sqrt{2+\sqrt{3}}/2\]

OpenStudy (solomonzelman):

/2 is outside the sqrt

OpenStudy (solomonzelman):

-.97 is the answer based on my work/solution!